OpenStudy (anonymous):

Write the slope intercept equationfor the line that passes through (2,-4) and has slope -7/2

OpenStudy (anonymous):

\[y=-\frac{7}{2}x+3\]

OpenStudy (anonymous):

The point-gradient form of a line is\[y-y_1=m(x-x_1)\]Here, m=-7/2 and \[(x_1,y_1)=(2,-4)\]So, upon substitution,\[y+4=-\frac{7}{2}(x-2) \rightarrow y=-\frac{7}{2}x+3\]